OpenStudy (anonymous):

After getting a discount of 20%, Jim paid $200 for a bicycle. How much did the bicycle originally cost?

OpenStudy (anonymous):

Let the original cost be x. x - 20x/100 = 200 Now you can solve the equation and get the answer.

OpenStudy (anonymous):

$240

OpenStudy (anonymous):

250$

OpenStudy (anonymous):

80x/100 = 200 x = 200 * 100/80 = 2000/8 = $ 250

OpenStudy (anonymous):

yeah, sorry it is $250!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw